OpenStudy (biohazard9064):

I really need help with this question. How are rational exponents related to integer exponents

OpenStudy (irishboy123):

if \( \large r =\frac{n}{d}\) then \( \large x^r = x^{\frac{n}{d}} = (x^{\frac{1}{d}})^{n} = (x^{n})^{\frac{1}{d}} = \sqrt[d]{x^n}\)

OpenStudy (biohazard9064):

ok

OpenStudy (biohazard9064):

is that all there is to it?

OpenStudy (irishboy123):

yes